OpenStudy (anonymous):

If f(x) = x^2e^-x , then the graph of f is decreasing for all x such that ? ...And there are answer choices to choose from. I just dont know how to set it up.

OpenStudy (across):

\[f(x)=x^2e^{-x}\]\[f'(x)=2xe^{-x}-x^2e^{-x}=xe^{-x}(2-x)\]

OpenStudy (anonymous):

So, would that mean x<0 and x >2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks